# SwiftTry: Fast and Consistent Video Virtual Try-On with Diffusion Models

### v1

@@ Abstract
Given an input video of a person and a new garment, the objective of this paper is to synthesize a new video where the person is wearing the specified garment while maintaining spatiotemporal consistency. While significant advances have been made in image-based virtual try-ons, extending these successes to video often results in frame-to-frame inconsistencies. Some approaches have attempted to address this by increasing the overlap of frames across multiple video chunks, but this comes at a steep computational cost due to the repeated processing of the same frames, especially for long video sequence. To address these challenges, we reconceptualize video virtual try-on as a conditional video inpainting task, with garments serving as input conditions. Specifically, our approach enhances image diffusion models by incorporating temporal attention layers to improve temporal coherence. To reduce computational overhead, we introduce ShiftCaching, a novel technique that maintains temporal consistency while minimizing redundant computations. Furthermore, we introduce the TikTokDress dataset, a new video try-on dataset featuring more complex backgrounds, challenging movements, and higher resolution compared to existing public datasets. Extensive experiments show that our approach outperforms current baselines, particularly in terms of video consistency and inference speed. Data and code are available at https://github.com/VinAIResearch/swift-try.
@@ Introduction
Figure 1: Results of our SwiftTry compared with those of ViViD (Fang et al. 2024), the previous method for video try-on. Our method maintains detail and consistency in the texture of the garments and runs over 60% faster.
@@
Video virtual try-on is an emerging research area (Chen et al. 2021; Rogge et al. 2014; Pumarola et al. 2019; Dong et al. 2019b; Kuppa et al. 2021; Zhong et al. 2021; Jiang et al. 2022; He et al. 2024; Xu et al. 2024b; Fang et al. 2024; Zheng et al. 2024) with significant potential in fashion and e-commerce. The ability to realistically visualize how a garment looks on a person in a video could transform online shopping. However, despite recent progress in image-based virtual try-on (He, Song, and Xiang 2022; Choi et al. 2021; Lee et al. 2022; Xie et al. 2023; Zhu et al. 2023; Kim et al. 2023), extending these capabilities to video remains challenging due to the need for spatiotemporal consistency and the high computational costs of processing long sequences.
@@
A significant challenge in video virtual try-on is balancing the need for temporal coherence with the computational demands of processing long video sequences. Previous methods (Xu et al. 2024b; He et al. 2024; Fang et al. 2024) often struggle with temporal inconsistencies, leading to visual artifacts and flickering between frames, which undermines the realism of the virtual try-on experience. Additionally, the high computational cost of rendering high-quality results over extended sequences hampers the practicality of these approaches for real-world applications.
@@
Another challenge is lacking an evaluation dataset. The first public video try-on dataset VVT (Dong et al. 2019b) only covers basic pattern garments, form-fitting T-shirts, uniform backgrounds, static camera angles, and repetitive human motions. Recently, ViViD (Fang et al. 2024) released the first practical dataset for video virtual try-on; however, it struggles to handle in-the-wild scenarios, such as complex movements and diverse backgrounds, making it challenging to meet the demands of practical applications. Moreover, one reason for the poor quality video try-on result is primarily due to masks extracted using human parsing segmentation (Li et al. 2020) applied on each frame of the video.
@@
In this paper, we tackle these challenges by introducing two key contributions. First, we present a new high-quality dataset named TikTokDress consisting of 817 videos which is specifically designed for training and evaluating video virtual try-on models. This dataset features realistic scenes, diverse garment types, and complex movements, providing a robust foundation for advancing research in this field. Second, we introduce a novel video virtual try-on framework named SwiftTry, as shown in Fig. 1 which significantly reduces the computational cost of processing long video sequences while maintaining temporal consistency. Our framework is inspired by state-of-the-art diffusion-based image virtual try-on methods (Kim et al. 2023; Xu et al. 2024a; Choi et al. 2024) and incorporates temporal attention in the UNet architecture to train on video try-on data. During inference, we introduce a new technique called ShiftCaching, which ensures temporal coherence and smooth transitions between video clips while also reducing redundant computation compared to previous methods. Extensive experimental results indicate that our proposed SwiftTry framework, incorporating the aforementioned techniques, substantially surpasses other video virtual try-on methods in performance.
@@
  • We propose a new technique for video inference named ShiftCaching, which can ensure temporal smoothness between video clips and reduce redundant computation.
@@
Figure 2: Overview of Stage 2 of our SwiftTry framework. Stage 1 is similar except the input is single image frame and does not have temporal attention layers. Given an input video and a garment image, our method first extracts the masked video, corresponding masks, and pose sequence. The masked video is encoded into latent space by the VAE Encoder, which is then concatenated with noise, masks, and pose features before being processed by the Main U-Net. To inpaint the garment during the denoising process, we use a Garment U-Net and a CLIP encoder to extract both low- and high-level garment features. These features are integrated into the Main U-Net through spatial and cross-attention mechanisms.
@@
Recently, diffusion models have markedly enhanced the realism of images in generative tasks, leading to their increasing adoption in virtual try-on research. For instance, TryOnDiffusion (Zhu et al. 2023) presents a virtual try-on method utilizing two U-Nets, but it requires a large dataset of image pairs depicting the same person in various poses, which can be difficult to acquire. StableVITON (Kim et al. 2023) conditions the garment in a ControlNet (Zhang, Rao, and Agrawala 2023)-style using a zero cross-attention block, while IDM-VTON (Choi et al. 2024) proposes GarmentNet to encode low-level features combined with high-level semantic features extracted via IP-Adapter (Ye et al. 2023). Despite these advancements, extending these existing image virtual try-on methods for video often results in significant inter-frame inconsistency and flickering, which adversely affects the overall quality of the generated results.
Video Virtual Try-On. Several efforts have been made to develop virtual try-on for videos. FW-GAN (Dong et al. 2019b) integrates an optical flow prediction module from Video2Video (Wang et al. 2018b) to warp preceding frames to the current frame, enabling the synthesis of temporally coherent subsequent frames. MV-TON (Zhong et al. 2021) introduces a memory refinement module that retains and refines features from previous frames. ClothFormer (Jiang et al. 2022) utilizes a vision transformer in its try-on generator to minimize blurriness and temporal artifacts, and it features an innovative warping module that combines TPS-based and appearance-based methods to address issues like incorrect warping due to occlusions. Among diffusion-based methods, Tunnel Try-On (Xu et al. 2024b) is the first to apply diffusion models for video virtual try-on, effectively handling camera movement and maintaining consistency, although its demo videos are limited to only a few seconds long. ViViD (Fang et al. 2024) introduced a large-scale video try-on dataset with multiple categories, but it remains limited by simple backgrounds and movements, which constrains its ability to maintain long-term consistency and coherence. In this paper, we propose a novel technique that ensures temporal smoothness and coherence across video clips, complemented by a caching technique (Ma, Fang, and Wang 2024) that reduces redundant computations during long video inference.
@@ Methods
Problem Statement: Given a source video $V=\{I_{1},I_{2},\dots,I_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of a person and a garment image $g\in\mathbb{R}^{3\times H\times W}$, where $N$, $H$, and $W$ represent the video length, frame height, and frame width, respectively, our goal is to synthesize an target video $\hat{V}=\{\hat{I}_{1},\hat{I}_{2},\dots,\hat{I}_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of the person wearing the garment, while preserving the motion of the person, the background in $V$, and the color and texture of $g$.
It is important to note that collecting both source and target videos of the same person with identical motion and gestures, differing only in the garment, is extremely challenging. As a result, most video try-on approaches adopt a self-supervised training method, where only a single video is used, and the garment regions are masked. The model is then trained to inpaint the masked regions using guidance from the garment image.
@@
Figure 3: Illustration of fully and partially computed frames with a chunk size $N=8$ and shift $\Delta=4$. In the partially computed chunk, one half uses cached features from timestep $t+2$, and the other half uses features from $t+1$.
@@
Our approach consists of two stages: first, training a diffusion-based image try-on model, and then extending it to work with video data by incorporating temporal attention into every block of the Main UNet.
In the first stage, inspired by StableVITON (Kim et al. 2023), we design diffusion-based image try-on model with two submodules: the Garment UNet and the Main UNet, as illustrated in Fig. 2. The Main UNet is a modified inpainting model, initialized with the pretrained weights from Stable Diffusion (Rombach et al. 2022). It takes as input four channels of latent noise, four channels of latent representations of the masked image (i.e., the person image with the clothes region masked), and one channel for the binary mask representing the inpainting region. To further enhance the generation, we add the pose skeleton as an additional control, represented by a pose map rendered from DW-Pose (Yang et al. 2023). This results in a 13-channel input fed into the Main UNet, which predicts the cleaned latent over $T$ timesteps. Finally, the cleaned latent is passed through a decoder to produce the output image. The Garment UNet has a similar architecture to the Main UNet but only takes the garment image as its input, rather than the multiple channels used in the Main UNet. This module is designed to extract both detailed and high-level features from the garment, guiding the Main UNet to accurately replicate the garment’s appearance through Reference Attention. Specifically, we follow the Reference Attention from AnimateAnyone (Hu et al. 2023), replicating the Garment Unet’s feature maps along the temporal dimension and concatenating them with the Main Unet’s feature maps along the spatial dimension before undergoing Unet’s self-attention.
We use the VITON-HD dataset (Choi et al. 2021) to train our network during this stage.
@@
Figure 4: Comparison of a fully computed frame and a partially computed frame. The partially computed frame uses Masked Temporal Attention instead of standard Temporal Attention to address mismatches in cached features.
In the second stage of designing our diffusion-based video try-on model, we ensure temporal consistency across video frames by adapting the Main UNet from the image try-on model. This adaptation involves converting its 2D layers into pseudo-3D layers (Guo et al. 2023; Wu et al. 2023; Zhou et al. 2022) and adding a temporal attention layer after the Spatial and Cross-Attention layers to capture temporal correlations between frames. The architecture of the modified UNet blocks is illustrated in Fig. 2. In the temporal attention layer, the features are reshaped into the shape of $(H\times W)\times N\times C$, where $C$ is the number of feature channels and $H\times W$ is the batch-size dimension, to compute self-attention along the temporal dimension. This allows a location in the latent space of frame $t$ to interact with the same location in other frames within a chunk of $N$ frames. This design is highly efficient as it avoids the expense of full 3D attention by factorizing it into two consecutive steps: spatial attention (2D) and temporal attention (1D). This approach allows a location in one frame to exchange information with every location in all other frames. Additionally, we incorporate sinusoidal positional encoding to help the model recognize the position of each frame in the video, following (Guo et al. 2023). In this stage, we train only the temporal attention layer while keeping the other layers unchanged, using a video dataset.
@@
Due to memory constraints, current video diffusion-based virtual try-on methods can only generate video chunks of 16 frames at a time. Previous approaches (Fang et al. 2024; He et al. 2024; Xu et al. 2024b) use a temporal aggregation technique (Tseng, Castellon, and Liu 2023; Xu et al. 2023) to stitch overlapping video clips into longer sequences. In this process, the long video is divided into overlapping clips with an overlap size $S$, typically set to $N/2$ or $N/4$. At each denoising timestep $t$, the overlapping noise predictions are merged using a simple averaging technique. However, this method involves a trade-off: a smaller overlap size, such as $S=4$, can cause temporal flickering and texture artifacts, while a larger overlap size, such as $S=15$, improves consistency but greatly slows down the process as shown in Tab. 1.
@@
To further accelerate the inference process, we can skip a random chunk to reduce redundant computation during denoising. However, dropping chunks without adjustment can lead to abrupt changes in noise levels in the final results. Following (Ma, Fang, and Wang 2024), which notes that adjacent denoising steps share significant similarities in high-level features, we instead perform partial computations on the Main U-Net. Specifically, we use a cache to copy the latest features from the fully computed timestep $z_{t+1}$ (Red frame) and use these features to partially compute the current latent $z_{t}$ (White frame), bypassing the deeper blocks of the UNet, as illustrated in Fig. 4.
When performing partial computations on a chunk, the cached features typically include a first half from timestep $t+2$ and a second half from timestep $t+1$, which can cause mismatches between the two halves. To address this, we introduce a Masked Temporal Attention mechanism. This mechanism incorporates a special mask with size $N\times N$ during the softmax attention calculation to set the attention matrix values to 0, preventing information transfer from less accurate features (timestep $t+2$) to more accurate features (timestep $t+1$), while allowing transfer from good features to bad features. This mechanism ensures both smoothness and high quality in the partially computed cells.
@@
Male Female White Yellow Black Bottom Top Center Near Far Move Stay Dynamic Static
@@
Table 2: Data statistics illustrating our dataset’s diversity and complexity. The table breaks down attributes such as gender, skin tone, camera positions, distances (Near/Far), and actions (Move/Stay), which indicate whether the actor is moving or stationary.
@@
Figure 6: SAM 2 (Ravi et al. 2024) failures due to sensitivity to prompts (green for positive, red for negative), with manual corrections needed for challenging areas (green rectangles).
Public datasets for single-image virtual try-ons, such as VITON-HD and DressCode, often suffer from simple backgrounds and limited human poses, despite offering high-quality images. These datasets are also restricted to single-image scenarios. Similarly, the VVT dataset, a standard for video virtual try-on, has notable drawbacks, including uniform movements, white backgrounds, and low resolution ($256\times 192$), making it unsuitable for real-world applications, particularly in the short-video industry where higher resolution is crucial. Given that real-world videos are typically recorded on mobile phones, which introduces variations in background, camera position, and lighting, there’s a need for a more robust dataset. To address these gaps, we introduce TikTokDress, a high-resolution video virtual try-on dataset that includes complex backgrounds, diverse movements, and gender representation. Each video is paired with the corresponding garment and is annotated with detailed human poses and precise binary cloth masks, enhancing its value for real-world applications.
@@
Our dataset consists of short TikTok clips (10-30 seconds) featuring various dance routines, as shown in Fig. 5. We expanded the TikTok Dataset (Jafarian and Park 2021) by adding videos to increase diversity in backgrounds, skin tones, and clothing styles, resulting in 817 video-garment pairs. Videos with excessive motion blur or low-quality garments were excluded. To ensure accurate garment matching, we manually selected high-quality matches from fashion retail websites. The dataset includes over 270,000 RGB frames extracted at 30 frames per second. We also calculated 2D keypoints and dense pose information using DWPose (Yang et al. 2023) and DensePose (Güler, Neverova, and Kokkinos 2018). Dataset statistics are summarized in Tab. 2, highlighting diversity in gender, skin tone, and camera positions.
Creating high-quality garment masks for each video was challenging due to the need for precise segmentation in every frame. We used SAM 2 (Ravi et al. 2024) to extract masks for both clothing and arms, but its sensitivity to prompt points and specific frames (see Fig. 6-(a)) required an additional solution. We developed an algorithm (described in the Supplemental Material) to automate the selection of optimal frames and prompts, improving efficiency across various scenarios. Despite these advancements, complex or unusual garments still needed manual intervention, which we detail with examples in the Supplementary Material. This meticulous process was essential for ensuring the dataset’s high quality and reliability.

### v2

@@ Abstract
Given an input video of a person and a new garment, the objective of this paper is to synthesize a new video where the person is wearing the specified garment while maintaining spatiotemporal consistency. Although significant advances have been made in image-based virtual try-on, extending these successes to video often leads to frame-to-frame inconsistencies. Some approaches have attempted to address this by increasing the overlap of frames across multiple video chunks, but this comes at a steep computational cost due to the repeated processing of the same frames, especially for long video sequences. To tackle these challenges, we reconceptualize video virtual try-on as a conditional video inpainting task, with garments serving as input conditions. Specifically, our approach enhances image diffusion models by incorporating temporal attention layers to improve temporal coherence. To reduce computational overhead, we propose ShiftCaching, a novel technique that maintains temporal consistency while minimizing redundant computations. Furthermore, we introduce the TikTokDress dataset, a new video try-on dataset featuring more complex backgrounds, challenging movements, and higher resolution compared to existing public datasets. Extensive experiments demonstrate that our approach outperforms current baselines, particularly in terms of video consistency and inference speed. The project page is available at https://swift-try.github.io/.
@@ Introduction
Figure 1: Results of our SwiftTry compared to those of ViViD (Fang et al. 2024), a previous method for video try-on. Our method preserves garment texture detail and consistency while achieving over 60% faster runtime.
@@
Video virtual try-on is an emerging research area (Chen et al. 2021; Rogge et al. 2014; Pumarola et al. 2019; Dong et al. 2019b; Kuppa et al. 2021; Zhong et al. 2021; Jiang et al. 2022; He et al. 2024; Xu et al. 2024b; Fang et al. 2024; Zheng et al. 2024) with significant potential in fashion and e-commerce. The ability to realistically visualize how a garment appears on a person in a video could transform online shopping. However, despite recent advances in image-based virtual try-on (He, Song, and Xiang 2022; Choi et al. 2021; Lee et al. 2022; Xie et al. 2023; Zhu et al. 2023; Kim et al. 2023), extending these capabilities to video remains challenging due to the need for spatiotemporal consistency and the high computational costs of processing long sequences.
@@
A significant challenge in video virtual try-on is balancing the need for temporal coherence with the computational demands of processing long video sequences. Previous methods (Xu et al. 2024b; He et al. 2024; Fang et al. 2024) often struggle with temporal inconsistencies, resulting in visual artifacts and flickering between frames, which undermines the realism of the virtual try-on experience. Additionally, the high computational cost of rendering high-quality results over extended sequences limits the practicality of these approaches in real-world applications.
@@
Another challenge is the lack of an adequate evaluation dataset. The first public video try-on dataset, VVT (Dong et al. 2019b), only includes basic pattern garments, form-fitting T-shirts, uniform backgrounds, static camera angles, and repetitive human motions. More recently, ViViD (Fang et al. 2024) introduced the first practical dataset for video virtual try-on. However, it struggles to handle in-the-wild scenarios, such as complex movements and diverse backgrounds, making it difficult to meet the demands of real-world applications. Moreover, the poor quality of video try-on results can often be attributed to the inaccurate masks extracted using human parsing segmentation (Li et al. 2020), which are applied to each frame of the video.
@@
In this paper, we address these challenges with two key contributions. First, we introduce a new high-quality dataset, named TikTokDress, consisting of 817 videos specifically designed for training and evaluating video virtual try-on models. This dataset features realistic scenes, diverse garment types, and complex movements, providing a robust foundation for advancing research in this field. Second, we propose a novel video virtual try-on framework named SwiftTry, as illustrated in Fig. 1, which significantly reduces the computational cost of processing long video sequences while maintaining temporal consistency.
Our framework is inspired by state-of-the-art diffusion-based image virtual try-on methods (Kim et al. 2023; Xu et al. 2024a; Choi et al. 2024) and incorporates temporal attention within the UNet architecture to train on video try-on data. During inference, we introduce a new technique called ShiftCaching, which ensures temporal coherence and smooth transitions between video chunks while minimizing redundant computation compared to previous methods. Extensive experimental results demonstrate that our proposed SwiftTry framework, leveraging these techniques, significantly outperforms existing video virtual try-on methods in both accuracy and efficiency.
@@
  • We propose a new technique for video inference named ShiftCaching, which can ensure temporal smoothness between video chunks and reduce redundant computation.
@@
Figure 2: Overview of Stage 2 of our SwiftTry framework (Note that stage 1 is similar, except the input is a single image frame, and it does not include temporal attention layers). Given an input video and a garment image, our method first extracts the masked video, corresponding masks, and pose sequence. The masked video is encoded into the latent space by the VAE Encoder, which is then concatenated with noise, masks, and pose features before being processed by the Main U-Net. To inpaint the garment during the denoising process, we use a Garment U-Net and a CLIP encoder to extract both low- and high-level garment features. These features are integrated into the Main U-Net through spatial and cross-attention mechanisms.
@@
Recently, diffusion models have markedly enhanced the realism of images in generative tasks, leading to their increasing adoption in virtual try-on research. For instance, TryOnDiffusion (Zhu et al. 2023) presents a virtual try-on method utilizing two U-Nets, but it requires a large dataset of image pairs of the same person in different poses, which can be difficult to acquire. StableVITON (Kim et al. 2023) conditions the garment in a ControlNet (Zhang, Rao, and Agrawala 2023)-style using a zero cross-attention block, while IDM-VTON (Choi et al. 2024) proposes GarmentNet to encode low-level features combined with high-level semantic features extracted via IP-Adapter (Ye et al. 2023). Despite these advancements, extending these existing image virtual try-on methods for video often results in significant inter-frame inconsistency and flickering, which adversely affects the overall quality of the generated results.
Video Virtual Try-On. Several efforts have been made to develop virtual try-on systems for videos. FW-GAN (Dong et al. 2019b) incorporates an optical flow prediction module from Video2Video (Wang et al. 2018b) to warp preceding frames to the current frame, enabling the synthesis of temporally coherent subsequent frames. MV-TON (Zhong et al. 2021) introduces a memory refinement module that retains and refines features from previous frames. ClothFormer (Jiang et al. 2022) employs a vision transformer in its try-on generator to minimize blurriness and temporal artifacts. It also features an innovative warping module that combines TPS-based and appearance-based methods to address challenges such as incorrect warping caused by occlusions. Among diffusion-based methods, Tunnel Try-On (Xu et al. 2024b) is the first to apply diffusion models for video virtual try-on, effectively handling camera movement and maintaining consistency. However, its demo videos are limited to only a few seconds in length. ViViD (Fang et al. 2024) introduced a large-scale video try-on dataset with multiple categories, but it remains limited by simple backgrounds and movements, which constrain its ability to ensure long-term consistency and coherence. In this paper, we propose a novel approach that establishes temporal smoothness and coherence across video chunks. Additionally, we integrate a caching technique (Ma, Fang, and Wang 2024) to reduce redundant computations during long video inference, significantly improving efficiency.
@@ Methods
Problem Statement: Given a source video $V=\{I_{1},I_{2},\dots,I_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of a person and a garment image $g\in\mathbb{R}^{3\times H\times W}$, where $N$, $H$, and $W$ represent the video length, frame height, and frame width, respectively, our goal is to synthesize a target video $\hat{V}=\{\hat{I}_{1},\hat{I}_{2},\dots,\hat{I}_{N}\}\in\mathbb{R}^{N\times 3\times H\times W}$ of the person wearing the garment, while preserving the motion of the person, the background in $V$, and the color and texture of $g$.
It is important to note that collecting both source and target videos of the same person with identical motions and gestures, differing only in the garment, is extremely challenging. As a result, most video try-on approaches adopt a self-supervised training method, where only a single video is used, and the garment regions are masked. The model is then trained to inpaint the masked regions using guidance from the garment image.
@@
Figure 3: Illustration of fully and partially computed frames with a chunk size of $N=8$ and a shift of $\Delta=4$. In the partially computed chunk, one half uses cached features from $t+2$, while the other half uses features from $t+1$.
@@
Our approach comprises two stages: first, training a diffusion-based image try-on model, and second, extending it to work with video data by incorporating temporal attention into every block of the Main UNet.
In the first stage, inspired by StableVITON (Kim et al. 2023), we design a diffusion-based image try-on model with two submodules: the Garment UNet and the Main UNet, as illustrated in Fig. 2. The Main UNet is a modified inpainting model initialized with pretrained weights from Stable Diffusion (Rombach et al. 2022). It takes as input four channels of latent noise, four channels of latent representations of the masked image (i.e., the person image with the clothing region masked), and one channel for the binary mask representing the inpainting region. To further enhance generation quality, we add the pose skeleton as an additional control, represented by a pose map rendered from DW-Pose (Yang et al. 2023). This results in a 13-channel input, which is fed into the Main UNet to predict the cleaned latent over $T$ timesteps. Finally, the cleaned latent is passed through a decoder to produce the output image.
The Garment UNet has a similar architecture to the Main UNet but only takes the garment image as its input, rather than the multiple channels used in the Main UNet. This module is designed to extract both detailed and high-level features from the garment, guiding the Main UNet to accurately replicate the garment’s appearance through Reference Attention. Specifically, we follow the Reference Attention mechanism from AnimateAnyone (Hu et al. 2023), replicating the Garment UNet’s feature maps along the temporal dimension and concatenating them with the Main UNet’s feature maps along the spatial dimension before applying the UNet’s self-attention. We use the VITON-HD dataset (Choi et al. 2021) to train our network during this stage.
@@
Figure 4: Comparison between a fully computed frame and a partially computed frame. The partially computed frame employs Masked Temporal Attention instead of standard Temporal Attention to resolve mismatches in cached features.
In the second stage, we modify the Main UNet from the image try-on model to ensure temporal consistency across video frames. This adaptation involves converting its 2D layers into pseudo-3D layers (Guo et al. 2023; Wu et al. 2023; Zhou et al. 2022) and adding a temporal attention layer after the Spatial and Cross-Attention layers to capture temporal correlations between frames. The architecture of the modified UNet blocks is illustrated in Fig. 2. In the temporal attention layer, the features are reshaped into the shape of $(H\times W)\times N\times C$, where $C$ is the number of feature channels and $H\times W$ is the batch-size dimension, to compute self-attention along the temporal dimension. This allows a location in the latent space of frame $t$ to interact with the same location in other frames within a chunk of $N$ frames. This design is highly efficient as it avoids the expense of full 3D attention by factorizing it into two consecutive steps: spatial attention (2D) and temporal attention (1D). This approach allows a location in one frame to exchange information with every location in all other frames. Additionally, we incorporate sinusoidal positional encoding to help the model recognize the position of each frame in the video, following (Guo et al. 2023). In this stage, we train only the temporal attention layer while keeping the other layers unchanged, using a video dataset.
@@
Due to memory constraints, current video diffusion-based virtual try-on methods can only generate video chunks of 16 frames at a time. Previous approaches (Fang et al. 2024; He et al. 2024; Xu et al. 2024b) use a temporal aggregation technique (Tseng, Castellon, and Liu 2023; Xu et al. 2023) to stitch overlapping video chunks into longer sequences. In this process, the long video is divided into overlapping chunks with an overlap size $S$, typically set to $N/2$ or $N/4$. At each denoising timestep $t$, the overlapping noise predictions are merged using a simple averaging technique. However, this method involves a trade-off: a smaller overlap size, such as $S=4$, can cause temporal flickering and texture artifacts, while a larger overlap size, such as $S=15$, improves consistency but greatly slows down the process as shown in Tab. 1.
@@
To further accelerate the inference process, we can skip a random chunk to reduce redundant computation during denoising. However, naively dropping chunks without adjustment can lead to abrupt changes in noise levels in the final results. Following (Ma, Fang, and Wang 2024), which notes that adjacent denoising steps share significant similarities in high-level features, we instead perform partial computations on the Main U-Net. Specifically, we use a cache to copy the latest features from the fully computed timestep $z_{t+1}$ (Red frame) and use these features to partially compute the current latent $z_{t}$ (White frame), bypassing the deeper blocks of the UNet, as illustrated in Fig. 4.
When performing partial computations on a chunk, the cached features typically include the first half from timestep $t+2$ and the second half from timestep $t+1$, which can lead to mismatches between the two halves. To address this, we introduce a Masked Temporal Attention mechanism. This mechanism applies a special mask of size $N\times N$ during the softmax attention calculation to set specific values in the attention matrix to 00. This prevents the transfer of information from less accurate features (timestep $t+2$) to more accurate features (timestep $t+1$), while allowing transfer from good features to bad features. This approach ensures both smoothness and high quality in the partially computed cells.
@@
Male Female White Asian Black Bottom Top Center Near Far Move Stay Dynamic Static
@@
Table 2: Data statistics highlighting the diversity and complexity of our dataset. The table provides a breakdown of attributes such as gender, skin tone, camera positions, distances (Near/Far), and actions (Move/Stay), indicating whether the actor is moving or stationary
@@
Figure 6: SAM 2 failures due to its sensitivity to prompts (green for positive, red for negative), requiring manual corrections for challenging areas (green rectangles).
Public datasets for single-image virtual try-ons, such as VITON-HD and DressCode, often suffer from simple backgrounds and limited human poses, despite offering high-quality images. These datasets are also restricted to single-image scenarios. Similarly, the VVT dataset, a standard for video virtual try-on, has notable drawbacks, including uniform movements, white backgrounds, and low resolution ($256\times 192$), making it unsuitable for real-world applications, particularly in the short-video industry where higher resolution is crucial. Since real-world videos are typically recorded on mobile phones, which introduce variations in background, camera position, and lighting, there is a pressing need for a more robust dataset. To address these shortcomings, we introduce TikTokDress, a high-resolution video virtual try-on dataset that includes complex backgrounds, diverse movements, and balanced gender representation. Each video is paired with its corresponding garment and annotated with detailed human poses and precise binary cloth masks, enhancing its utility for real-world applications.
@@
Our dataset consists of short TikTok clips (10–30 seconds) showcasing various dance routines, as illustrated in Fig. 5. We expanded the TikTok Dataset (Jafarian and Park 2021) by adding videos to enhance diversity in backgrounds, skin tones, and clothing styles, resulting in 817 video-garment pairs. Videos with excessive motion blur or low-quality garments were excluded. To ensure accurate garment matching, we manually curated high-quality matches from fashion retail websites. The dataset includes over 270,000 RGB frames extracted at 30 frames per second. Additionally, we computed 2D keypoints and dense pose information using DWPose (Yang et al. 2023) and DensePose (Güler, Neverova, and Kokkinos 2018). Dataset statistics, summarized in Tab. 2, highlight its diversity in gender, skin tone, and camera positions.
Creating high-quality garment masks for each video posed a significant challenge due to the need for precise segmentation in every frame. We used SAM 2 (Ravi et al. 2024) to extract masks for both clothing and arms. However, its sensitivity to prompt points and specific frames (see Fig. 6-(a)) necessitated an additional solution. To improve efficiency, we developed an algorithm (detailed in the Supplemental Material) for optimal frame and prompt selection. Complex garments still required manual refinement, as shown in the Supplementary Material. This meticulous process was essential for ensuring the dataset’s high quality and reliability.